Simlify. (x-y/x+y)+(x+y/x-y)
x2-y+xy/x+x3-xy2

Found 2 solutions by fractalier, Kathy:
Answer by fractalier(2101)

We can use the criss-cross method for this one...
(x-y/x+y)+(x+y/x-y) =
[(x - y)^2 + (x + y)^2] / (x^2 - y^2) =
[x^2 - 2xy + y^2 + x^2 + 2xy + y^2) / (x^2 - y^2) =
2(x^2 + y^2) / (x^2 - y^2)
The criss-cross method is
A/B ± C/D = (AD ± BD) / BD

Answer by Kathy(4)

x2-y+xy x3-xy2 or x2-y+xy/2x3-xy2
----
2